# FeDaL: Federated Dataset Learning for Time Series Foundation Models

### v1

@@ 4.3 Federated Scaling Behaviors
While prior work has examined scaling laws of centralized TSFMs, we investigate how federated pretraining scales with respect to (i) data size, (ii) number of clients, and (iii) client participation rate. Specifically, we vary: (i) data size from 40B to $\sim$231B with fixed 174 clients; (ii) client count from 30 to 174 under fixed total data; and (iii) participation rate from 10% to 100%. As shown in Figure˜6, larger data consistently improves performance, more clients yield better representations (even under fixed total data), and higher participation enhances aggregation and mitigates drift. These results indicate that federated TSFM pretraining benefits from scaling in data and client diversity, emphasizing coverage and participation over model size for improved generalization.
Figure 6: Scaling behaviors across tasks. Color codes: Blue – Avg. full/few-shot forecasting, Orange – Zero-shot forecasting, Green – Imputation, Red – Classification, Purple – Anomaly Detection. Y-axis for forecasting/imputation is inverted (lower is better). Full details and plots in Section˜C.5.
@@ 5 Conclusion
We propose Federated Dataset Learning (FeDaL), a brand-new framework for pretraining TSFMs under federated setting. By introducing Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), Our FeDaL effectively mitigates domain bias caused by heterogeneous time series datasets, enabling the learning of domain-invariant temporal representations. Experiments on eight downstream tasks confirm FeDaL’s strong cross-domain generalization and superior performance. We also provide the first federated scaling analysis, offering empirical guidance for future TSFM design.
@@
  • More Related Work (Appendix˜A): A comprehensive review of relevant literature, covering time series foundation models, FL for heterogeneous data, and foundation models.
  • Implementation Details (Appendix˜B): Detailed descriptions of the techniques employed, training configurations, experimental setup—including benchmark procedures, data processing pipelines, and baseline methods.
  • Full Results (Appendix˜C): Complete presentation of all results discussed in the main text.
@@ Appendix A More Related Work
Time Series Foundation Models
@@
Pre-trained models with scalability can evolve into foundation models, characterized by increasing model capacity and pre-training scale, enabling them to address diverse data. Large Language Models (LLMs) have demonstrated advanced capabilities, including in-context learning and few/zero-shot learning [34]. However, due to the heterogeneity of time series data, the development of GPT-style time series foundation models (TSFMs) has been limited. Research on TSFMs is in its early stages, with existing efforts categorized into two groups, one of which is LLM-empowered time series analysis. For instance, FPT [82] utilizes GPT-2 as a representation extractor for time series, fine-tuning it on downstream datasets and tasks. LLM4TS [7] encodes time series into numerical tokens for LLMs, demonstrating model scalability in forecasting tasks. Time-LLM [34] explores prompting techniques to enhance prediction, inspiring cross-modality representation ability through language templates. UniTime [46] introduces cross-domain techniques to fine-tune pre-trained LLMs, reducing prediction biases. Nevertheless, the performance of these LLM-based approaches is heavily dependent on the LLM backbone and cross-modal design. In contrast, our proposed model is pre-trained natively on time series, eliminating the need for extra modality alignment. Another category of models involves pre-training on large-scale time series datasets. For instance, ForecastFPN [23] is trained on synthetic series to achieve zero-shot forecasting. CloudOps [63] employs masked modeling on Transformers for domain-specific forecasting. TimeGPT-1 [27] introduces the first commercial API for zero-shot forecasting, marking a milestone in the field. PreDcT [18], a decoder-only Transformer pre-trained on Google Trends, demonstrates notable zero-shot capabilities. Furthermore, MOMENT [28], Moirai [64], TimeFM [19], Chronos [2], and Timer [48] pre-train on ultra-large time series datasets and release corresponding datasets, showcasing excellent cross-task performance empirically. This suggests that training TSFMs is a resource-intensive endeavor.
Federated Learning for Heterogeneous Data
@@
Federated Learning (FL) [53] enables decentralized model training across distributed clients without sharing raw data. A key challenge in FL is statistical heterogeneity (non-IID data) [13], which often leads to performance degradation due to distributional shifts across clients. Early solutions reduce client-server divergence through alignment or regularization, but typically assume mild heterogeneity and shared representation spaces—assumptions that break down in highly diverse domains. To address this, Personalized Federated Learning (PFL) aims to tailor models to individual clients. Existing methods include regularization-based decomposition [30, 40], partial model sharing [42, 16], adaptive aggregation [76], and meta-learning strategies [25]. While effective in some settings, PFL generally targets image or text data and assumes transferable low-level features [9], which do not generalize well to time series due to deeper cross-domain disparities in resolution, semantics, and physical context. Moreover, PFL prioritizes personalized models over unified generalization, making it ill-suited for time series foundation model (TSFM) pretraining, where cross-domain generalization is essential. To fill this gap, we propose Federated Dataset Learning (FeDaL)—a unified framework designed for TSFM training in decentralized, heterogeneous environments. By introducing Domain Bias Elimination (DBE) and Global Bias Elimination (GBE), FeDaL explicitly mitigates domain-specific biases at both client and server levels, enabling robust, scalable learning of generalized temporal representations.
Federated Learning for Foundation Models
FL is emerging as a powerful paradigm for training and adapting foundation models [70, 8, 10], especially in scenarios where centralizing large-scale, heterogeneous data is impractical due to computational or privacy constraints [9, 83, 32]. Unlike centralized strategies that require massive data aggregation and uniform model scaling, FL distributes the workload across clients, allowing each to contribute without sharing raw data. This decentralization not only reduces the computational burden but also preserves privacy while naturally accommodating data heterogeneity [83, 32]. Recent advances have explored FL for fine-tuning LLMs in personalized settings [49, 71, 32], demonstrating its potential to adapt pretrained models to diverse user data. In the time series domain, FL has also been used to fine-tune or train models for domain-specific tasks [44]. However, training foundation models for time series from scratch in federated settings presents new opportunities: FL offers a natural framework to leverage cross-domain data without mixing them, avoiding undesirable artifacts from centralized blending and enabling more principled treatment of domain-specific structures [9]. Despite these advantages, cross-domain bias—a key obstacle to generalization—has received little attention in federated TSFM training. Our proposed FeDaL specifically designed to address this gap. By explicitly mitigating client- and domain-level biases, FeDaL enables the training of unified TSFMs that capture high-level temporal representations transferable across domains and tasks.
@@ Appendix B Implementation Details
We employ the Transformer as the model backbone due to its excellent scalability. Inspired by the significant advancements in Decoder-only LLMs capable of iterative generation [27, 82], and recognizing the need for processing variable-length time series [48, 74], we adopt an auto-regressive approach for generative pre-training using standard Decoder-only Transformer architectures. The next-token prediction can be formulated as:
@@
$\mathcal{F}(\mathcal{C}_{\text{init}})=A+iP,$ (10)
where $A$ and $P$ denote the amplitude and phase components, respectively. We then perturb only the amplitude component with Gaussian noise:
$\tilde{\mathcal{C}}=\mathcal{F}^{-1}\left((A+\epsilon\cdot\mathcal{N}(0,1))+iP\right),$ (11)
where $\mathcal{F}^{-1}$ denotes the inverse Fourier transform, and $\epsilon$ controls the perturbation strength. This ensures that while the fine-grained raw sequence details are obfuscated, the high-level temporal semantics embedded in the phase remain intact. To ensure that $\mathcal{C}^{\prime}$ remains semantically faithful to the original data, we align its latent representation with that of the sampled mini-batch $\tilde{\mathcal{X}}$:
$\mathcal{L}_{\text{align}}({\mathcal{C}}^{\prime})=\frac{1}{K}\|\sum\nolimits_{c_{j}\in\mathcal{C}^{\prime}}f_{\theta}(c_{j})-\sum\nolimits_{x_{k}\in\tilde{\mathcal{X}}}f_{\theta}(x_{k})\|_{2}^{2}.$ (12)
The semantically aligned core-set $\mathcal{C}$, optimized from the perturbed $\mathcal{C}^{\prime}$ via Eq. 12, is then uploaded to the server for the server-side core-set tuning via the privacy-preserving manner.
@@
To further assess the semantic fidelity and privacy-preserving effect of our core-set pipeline, we visualize the t-SNE embeddings of three data subsets for several clients: (i) the original local mini-batch $\tilde{\mathcal{X}}$ used for core-set construction, (ii) the gradient-matched core-set $\mathcal{C}_{\text{init}}$, and (iii) the final perturbed and aligned core-set $\mathcal{C}^{\prime}$ sent to the server. As shown in Figure˜7, the initial core-set (green) exhibits tight overlap with the original batch (pink), confirming that $\mathcal{C}_{\text{init}}$ successfully approximates local gradients—but also indicating potential privacy risks if shared directly. In contrast, the perturbed and aligned core-set (blue) is visibly displaced, thanks to our Fourier-based amplitude perturbation. Nonetheless, $\mathcal{C}^{\prime}$ maintains global structure and cluster consistency with $\tilde{\mathcal{X}}$, highlighting that our representation alignment objective effectively preserves semantic integrity while masking raw data characteristics. These results validate our claim that the core-set tuning pipeline provides a strong utility-privacy trade-off in the server-side core-set tuning.
@@
Time series decomposition aims to separate a sequence into interpretable components, typically trend, seasonal, and residual parts. This allows models to better capture structured temporal patterns and isolate noise or biases. In our framework, we follow the decomposition strategy proposed in [68], applying it directly in the latent representation space rather than the raw input domain. Given a latent representation sequence $\mathbf{h}=f_{\theta^{b}}({\bm{X}}_{i}^{\text{masked}})\in\mathbb{R}^{L\times d}$, we decompose it into trend and seasonal components using a moving average operator with period $\tau$:
$\mathbf{h}^{\text{trend}}=\texttt{MA}_{\tau}(\mathbf{h}),\quad\mathbf{h}^{\text{season}}=\mathbf{h}-\mathbf{h}^{\text{trend}}.$ (13)
Here, $\texttt{MA}_{\tau}(\cdot)$ denotes a $\tau$-point moving average operator. The resulting $\mathbf{H}^{\text{trend}}$ captures low-frequency temporal dynamics, while $\mathbf{H}^{\text{season}}$ reflects higher-frequency but still structured periodic variations. We omit the residual component to reduce complexity and avoid modeling unstructured noise. To estimate persistent client-specific deviations, we apply temporal averaging over both components:
$\mathbf{b}_{i}^{\text{trend}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{trend}}),\quad\mathbf{b}_{i}^{\text{season}}=\texttt{Mean}(\mathbf{h}_{i}^{\text{season}}),$ (14)
which serve as compact and effective approximations of domain-specific biases accumulated from heterogeneous time series distributions.
Basic Training Configuration
For both federated representation learning and downstream adaptation, we set the batch size to 2048. The local training uses 10 epochs per round, with total global rounds of 200 for federated representation learning and 1000 for downstream adaptation. The initial learning rate is fixed at $1e^{-2}$. Additionally, we configure the core-set size as $K=1024$ for federated representation learning and $K=3072$ for downstream adaptation.
@@
We evaluate FeDaL across two major settings: Federated Representation Learning and Downstream Adaptation. For Federated Representation Learning, we use two large-scale cross-domain time series datasets: (1) UTSD [^1] https://huggingface.co/datasets/thuml/UTSD, containing over 1 billion time points across 7 domains (e.g., energy, environment, health, IoT, nature); and (2) CTSD [^2] https://forecastingdata.org/, sampled from the Monash Time Series Forecasting Repository, comprising 500 million time points from 6 domains including ergonomics, transportation, health, energy, nature, and web. For Downstream Adaptation, we pretrain on the large-scale LOTSA dataset [64]  [^3] https://huggingface.co/datasets/Salesforce/lotsa_data, which contains 231 billion observations spanning multiple domains. We evaluate the pretrained TSFM from FeDaL on four downstream tasks: long- and short-term forecasting, imputation, classification, and anomaly detection. For long-term forecasting and imputation, we follow the GPT4TS [82] protocol, using ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Illness datasets (details in Table˜8). For short-term forecasting, we adopt the M4 dataset, again following GPT4TS, covering various temporal granularities (details in Table˜9). For classification, we follow the TimesBERT [74] setup, using 10 representative datasets from UEA and 91 from UCR to cover diverse domains (see Table˜10). For anomaly detection, we adopt the FFTS [9] benchmark, evaluating on SMD, MSL, SMAP, SWaT, and PSM datasets (details in Table˜11). Any datasets overlapping with LOTSA have been excluded from the pretraining phase.

### v2

@@ 4.3 Federated Scaling Behaviors
While prior work has examined scaling laws of centralized TSFMs (Yao et al., 2024), we investigate how federated pretraining scales with respect to data size, number of clients, and client participation rate. Specifically, we vary: (i) data size from 40B to 231B with fixed 174 clients; (ii) client count from 30 to 174 under fixed total data; and (iii) participation rate from 10% to 100%. As shown in Figure 7, larger data consistently improves performance, more clients yield better representations (even under fixed total data), and higher participation enhances aggregation and mitigates drift. These results indicate that federated TSFM pretraining benefits from scaling in data and client diversity, emphasizing coverage and participation over model size for improved generalization.
Figure 7: Scaling behaviors across tasks. Color codes: Blue – Avg. full/few-shot forecasting, Orange – Zero-shot forecasting, Green – Imputation, Red – Classification, Purple – Anomaly Detection. Y-axis for forecasting/imputation is inverted (lower is better). Full details and plots in Section C.7.
4.4 Discussion on Larger Time Series Forecasting Models
[…]
where $\alpha=\beta=1$ by default. A higher IGC indicates a model training with better efficiency.
Table 8: Zero-shot long-term forecasting performance comparison with Time-MoE. Bold: the best, Underline: the second best. Full results are provided in Table 20.
Models FeDaL (Ours) Time-MoEbase Time-MoElarge Time-MoEultra
Metrics
MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1 96 0.347 0.381 0.357 0.381 0.350 0.382 0.349 0.379
192 0.398 0.410 0.384 0.404 0.388 0.412 0.395 0.413
336 0.425 0.452 0.411 0.434 0.411 0.430 0.447 0.453
720 0.457 0.469 0.449 0.477 0.427 0.455 0.457 0.462
Avg. 0.407 0.429 0.400 0.424 0.394 0.419 0.412 0.426
ETTh2 96 0.307 0.355 0.305 0.359 0.302 0.354 0.292 0.352
192 0.349 0.372 0.351 0.386 0.364 0.385 0.347 0.379
336 0.387 0.395 0.391 0.418 0.417 0.425 0.406 0.419
720 0.401 0.406 0.419 0.454 0.537 0.496 0.439 0.447
Avg. 0.361 0.382 0.366 0.404 0.405 0.415 0.371 0.399
ETTm1 96 0.289 0.346 0.338 0.368 0.309 0.557 0.281 0.341
192 0.317 0.369 0.353 0.388 0.346 0.381 0.305 0.358
336 0.370 0.420 0.381 0.413 0.373 0.408 0.369 0.395
720 0.464 0.426 0.504 0.493 0.475 0.477 0.469 0.472
Avg. 0.360 0.390 0.394 0.415 0.376 0.405 0.356 0.391
ETTm2 96 0.207 0.283 0.201 0.291 0.197 0.286 0.198 0.288
192 0.248 0.333 0.258 0.334 0.250 0.322 0.235 0.312
336 0.316 0.340 0.324 0.373 0.337 0.375 0.293 0.348
720 0.397 0.408 0.488 0.464 0.480 0.461 0.427 0.423
Avg. 0.292 0.341 0.317 0.365 0.316 0.361 0.288 0.344
Weather 96 0.159 0.212 0.160 0.214 0.159 0.213 0.157 0.211
192 0.217 0.264 0.210 0.260 0.215 0.266 0.208 0.256
336 0.285 0.312 0.274 0.309 0.291 0.322 0.255 0.290
720 0.359 0.348 0.418 0.405 0.415 0.400 0.405 0.397
Avg. 0.255 0.284 0.265 0.297 0.270 0.300 0.256 0.288
Average
0.335 0.370 0.343 0.382 0.355 0.387 0.342 0.369
Total Param.#
28.42 M 113 M 453 M 2.4 B
Training Data
$\sim$ 231B 300B 300B 300B
IGC
[…]
@@ 5 Conclusion
We introduced FeDaL, a federated learning framework for training TSFMs from scratch that explicitly targets dataset-level heterogeneity. By integrating DBE at the client side and GBE at server, FeDaL mitigates both local and global biases, enabling the learning of robust, domain-invariant temporal patterns. Extensive experiments across diverse time series analysis tasks (various regression and classification) demonstrate its superior cross-domain learning capabilities and show that FeDaL-trained models remain highly competitive against centralized TSFMs with far larger parameter scales.
Reproducibility Statement
We pretrain on publicly available datasets UTSD (Liu et al., 2024c), CTSD (Chen et al., 2025), and LOTSA (Woo et al., 2024). Experiments are conducted on widely used benchmarks, including TSLib (Wu et al., 2022a) (ETT, Weather, and M4 for forecasting; ETT and Weather for imputation; SMD, MSL, SMAP, SwaT, and PSM for anomaly detection), as well as the UEA (Bagnall et al., 2018) and UCR (Dau et al., 2019) archives. All datasets are publicly accessible. To ensure reproducibility, we detail training setups, model configurations, and hyperparameter choices in Appendix B and Table 9. Furthermore, the code implementation of our proposed FeDaL is released at https://github.com/shengchaochen82/FeDaL.
@@
  • More Related Work (Appendix A): A comprehensive review of relevant literature, covering time series foundation models, FL for heterogeneous data, and foundation models.
  • Implementation Details (Appendix B): Detailed descriptions of the techniques employed, training configurations, experimental setup, including benchmark procedures, data processing pipelines, and baseline methods.
  • Full Results (Appendix C): Complete presentation of all results discussed in the main text.
@@ Appendix A More Related Work
Pre-trained Time Series Models.
@@
Large Language Models (LLMs) have demonstrated strong capabilities such as zero-shot adaptation (Jin et al., 2023). Yet, due to the heterogeneity of time series, GPT-style TSFMs remain underexplored. Existing research can be grouped into two categories. The first leverages LLMs for time series analysis. For example, FPT (Zhou et al., 2023) uses GPT-2 as a representation extractor, while LLM4TS (Chang et al., 2023) encodes series into numerical tokens, showing scalability on forecasting tasks. Time-LLM (Jin et al., 2023) employs prompting techniques to enhance cross-modality reasoning, and UniTime (Liu et al., 2024b) introduces domain adaptation to reduce prediction bias. However, these approaches largely depend on the LLM backbone and cross-modal design. In contrast, our model is pre-trained natively on time series, avoiding extra modality alignment. The second line pre-trains directly on large-scale time series datasets. ForecastFPN (Dooley et al., 2024) trains on synthetic series for zero-shot forecasting, while CloudOps (Woo et al., 2023) applies masked modeling for domain-specific prediction. TimeGPT-1 (Garza and Mergenthaler-Canseco, 2023) represents the first commercial zero-shot API, and PreDcT (Das et al., 2023) shows strong results with a decoder-only Transformer. More recently, models such as MOMENT (Goswami et al., 2024), Moirai (Woo et al., 2024), TimeFM (Das et al., 2024), Chronos (Ansari et al., 2024), and Timer (Liu et al., 2024c) have been trained on ultra-large datasets and publicly released, achieving competitive cross-task performance. These efforts collectively highlight both the potential and the resource-intensive nature of training TSFMs.
Federated Learning for Heterogeneous Data.
@@
FL (McMahan et al., 2017; Feng and Chen, 2025; Chen and Shu, 2025; Yan et al., 2025; Chen et al., 2024a; 2023; b) enables decentralized model training without raw data sharing, but suffers from statistical heterogeneity (non-IID data), where distributional shifts across clients degrade performance. Early approaches reduce client–server divergence via alignment or regularization, yet typically assume mild heterogeneity and shared representation spaces, an assumptions that collapse in highly diverse domains. Personalized Federated Learning (PFL) extends FL by tailoring models to individual clients through regularization (Hanzely et al., 2020; Li et al., 2021a; Chen et al., 2024c), partial parameter sharing (Li et al., 2021b; Collins et al., 2021), adaptive aggregation (Zhang et al., 2020), or meta-learning (Fallah et al., 2020). While effective for vision or text tasks where low-level features transfer well (Chen et al., 2025), PFL is less suitable for time series, where dataset-level disparities in resolution, semantics, and physical context are more severe. Moreover, PFL emphasizes personalization over unified generalization, conflicting with the goal of pretraining TSFMs, which require cross-domain invariance.
@@ Appendix B Implementation Details
We employ the Transformer as the model backbone due to its excellent scalability. Inspired by the significant advancements in Decoder-only LLMs capable of iterative generation (Garza and Mergenthaler-Canseco, 2023; Zhou et al., 2023), and recognizing the need for processing variable-length time series (Liu et al., 2024c; Zhang et al., 2025b), we adopt an auto-regressive approach for generative pre-training using standard Decoder-only Transformer architectures. The next-token prediction can be formulated as:
@@
$\hat{\mathcal{C}}=\mathcal{F}(\mathcal{C}_{\text{init}})=A\odot e^{iP},$ (11)
where $A$ and $P$ denote amplitude and phase, and $\odot$ is the elementwise product. We then add Gaussian noise only to the amplitude:
$\tilde{A}=A+\epsilon\,\Xi,\qquad\Xi\sim\mathcal{N}(0,I),$ (12)
and reconstruct the perturbed core-set by
$\mathcal{C}’=\mathcal{F}^{-1}\big(\tilde{A}\odot e^{iP}\big),$ (13)
where $\epsilon>0$ controls the perturbation strength and $\mathcal{F}^{-1}$ is the inverse DFT. To keep $\mathcal{C}$’ semantically faithful to the client data while remaining privacy-preserving, we align its latent representations to those of the sampled mini-batch $\tilde{\mathcal{X}}$ (both of size $K$) via
$\mathcal{L}_{\text{align}}(\mathcal{C}’)=\Big\|\frac{1}{K}\!\sum_{c\in\mathcal{C}’}\!f_{\theta}(c)\;-\;\frac{1}{K}\!\sum_{x\in\tilde{\mathcal{X}}}\!f_{\theta}(x)\Big\|_{2}^{\,2},$ (14)
where $f\theta(\cdot)$ denotes the encoder that outputs latent embeddings. The semantically aligned core-set $\mathcal{C}$, obtained by optimizing $\mathcal{L}_{\text{align}}$ w.r.t. $\mathcal{C}$’, is then uploaded to the server for core-set tuning.
@@
To further examine the semantic fidelity and privacy robustness of our core-set pipeline, we visualize the t-SNE embeddings of three data subsets for several clients: (i) the original local mini-batch $\tilde{\mathcal{X}}$ used for core-set construction, (ii) the gradient-matched core-set $\mathcal{C}{\text{init}}$, and (iii) the final perturbed and aligned core-set $\mathcal{C}$’ sent to the server. As shown in Figure 8, the initial core-set (green) largely overlaps with the original batch (pink), indicating that $\mathcal{C}{\text{init}}$ effectively approximates local gradients but may still retain fine-grained information if shared directly. After applying Fourier-based perturbation and semantic alignment, the refined core-set (blue) becomes more separated from the raw data distribution while preserving overall cluster structure. This suggests that our perturbation mechanism reduces exposure of fine-grained data characteristics while maintaining high-level semantic consistency required for server-side alignment. Overall, the results indicate that the proposed pipeline achieves a practical utility–privacy trade-off without claiming formal privacy guarantees. We note that our perturbation mechanism mitigates but does not eliminate potential information leakage, and formal differential privacy guarantees remain future work.
@@
Time series decomposition separates a sequence into interpretable components, typically trend, seasonal, and residual parts, enabling models to capture structured temporal patterns while isolating noise or biases. In our framework, we adopt the decomposition strategy from (Wu et al., 2021), but apply it directly in the latent representation space instead of the raw input domain. Given the latent representation $\mathbf{h}=f_{\theta^{b}}(\tilde{{\bm{X}}})\in\mathbb{R}^{L\times d}$, we decompose it into trend and seasonal parts using a $\tau$-point moving average operator:
$\mathbf{h}_{\text{t}}=\operatorname{MA}_{\tau}(\mathbf{h}),\quad\mathbf{h}_{\text{s}}=\mathbf{h}-\mathbf{h}_{\text{t}}.$ (15)
Here, $\mathbf{h}_{\text{t}}$ captures low-frequency temporal dynamics, while $\mathbf{h}_{\text{s}}$ represents higher-frequency structured variations. The residual component is omitted to avoid modeling unstructured noise. To estimate persistent dataset-specific deviations, we average each component over the temporal dimension:
$\mathbf{b}_{\text{t}}=\operatorname{Mean}(\mathbf{h}_{\text{t}}),\quad\mathbf{b}_{\text{s}}=\operatorname{Mean}(\mathbf{h}_{\text{s}}),$ (16)
where $\mathbf{b}_{\text{t}}$ and $\mathbf{b}_{\text{s}}$ are compact approximations of dataset-level biases arising from heterogeneous time series distributions.
Training Configuration
We evaluate FeDaL from two complementary perspectives: (1) its ability to learn generic representations in heterogeneous, distributed settings (referred to as Federated Representation Learning), and (2) the transferability of the learned TSFM to various downstream time series tasks (Downstream Adaptation). The detailed configurations for both tasks are summarized in Table 9. For federated representation learning, we use the UTSD and CTSD datasets to assess FeDaL under two forms of cross-domain heterogeneity: domain-mixed (DM) and domain-independent (DI). For downstream adaptation, we pre-train the Decoder-only Transformer via FeDaL on LOTSA, a large-scale cross-domain dataset comprising 231 billion time points, to ensure broad coverage of temporal patterns prior to adaptation.
Table 9: Detailed training configuration across various tasks.
Category Configuration Configuration
Federated Representation Learning Pretraining for TSFMs
Optimizer SGD SGD
Number of Clients 10 (UTSD) / 18 (CTSD) 174 (LOTSA)
Batch Size 2048 2048
Client Join Ratio ($\rho$) 70% 70%
Local Epochs 10 10
Global Rounds 200 1000
Learning Rate (Local Updating) 0.01 0.01
Core-set Generation Epochs 5 5
Learning Rate (Core-set Generation) 0.005 0.01
Server-side Tuning 5 15
Learning Rate (Server-side Tuning) 0.01 0.01
Input Sequence Length 1024 3072
Learning Scheduler StepLR StepLR
[…]
Smoother Factor ($\mu$) 0.2 0.1
$\alpha$ 0.7 0.7
$\beta$ 0.1 0.1
$\lambda$ (Alignment) 0.01 0.1
Core-set Size ($K$) 1024 ($1\%$) 3072 ($0.01\%$)
Decomposition Period ($\tau$) 4 8
Model Dimension 512 768
Number of Heads 4 8
Decoder Layers 4 8
Feedforward Dimension 512 768
Dropout 0.1 0.35
Activation Function GELU GELU
Masking Ratio 75% 50%
Patch Length 32 64
Stride 32 32
Evaluation Pipeline Masked Patches Reconstruction Downstream Adaption
Downstream Tasks N/A Long-term Forecasting (Full-/Zero-shot)
Short-term Forecasting
Classification
Anomaly Detection
@@
We evaluate FeDaL across two major settings: Federated Representation Learning and Downstream Adaptation. For Federated Representation Learning, we use two large-scale cross-domain time series datasets: (1) UTSD [^1] https://huggingface.co/datasets/thuml/UTSD, containing over 1 billion time points across 7 domains (e.g., energy, environment, health, IoT, nature); and (2) CTSD [^2] https://forecastingdata.org/, sampled from the Monash Time Series Forecasting Repository, comprising 500 million time points from 6 domains including ergonomics, transportation, health, energy, nature, and web. For Downstream Adaptation, we pretrain on the large-scale LOTSA dataset (Woo et al., 2024)  [^3] https://huggingface.co/datasets/Salesforce/lotsa_data, which contains 231 billion observations spanning multiple domains. We evaluate the pretrained TSFM from FeDaL on four downstream tasks: long- and short-term forecasting, imputation, classification, and anomaly detection. For long-term forecasting and imputation, we follow the GPT4TS (Zhou et al., 2023) protocol, using ETTh1, ETTh2, ETTm1, ETTm2, Weather, and Illness datasets (details in Table 11). For short-term forecasting, we adopt the M4 dataset, again following GPT4TS, covering various temporal granularities (details in Table 12). For classification, we follow the TimesBERT (Zhang et al., 2025b) setup, using 10 representative datasets from UEA and 91 from UCR to cover diverse domains (see Table 13). For anomaly detection, we adopt the FFTS (Chen et al., 2025) benchmark, evaluating on SMD, MSL, SMAP, SWaT, and PSM datasets (details in Table 14). Any datasets overlapping with LOTSA have been excluded from the pretraining phase.